OpenStudy (anonymous):

5 ________________ (6 - 4√7) - (8 - 5√7)

OpenStudy (anonymous):

\[\sqrt{7-2}\]

OpenStudy (anonymous):

\[6 - 4\sqrt{7} - 8 + 5\sqrt{7}\] \[= -2 + \sqrt{7}\] will be the denominator

OpenStudy (anonymous):

and doodle tech, I think only the 7 will be under the root.

OpenStudy (anonymous):

Yeah only the 7 would be. I messed up on typing the equation.

OpenStudy (anonymous):

(5/3)*(-2+sqrt(7)) To get this you must rationalize the denominator. Times both top and bottom by (-2-sqrt(7))